# Joint Beamforming Optimization and Mode Selection for RDARS-aided MIMO Systems

### v1

@@ II-A RDARS Architecture
A RDARS composed of $N$ elements is employed, with $a$ elements connected to the BS and operating in the “connection mode”, while the remaining $N\!-\!a$ passive elements operate in the “reflection mode” to reflect the incident signal to favorable directions. Since each element can be dynamically switched between the connection and reflection modes, it offers an extra DoF to achieve selection gains. Such transformation can be realized via a switching network (SW) as illustrated in Fig. 1. Flexible configurations are available for the number and locations of elements in the connection mode. In this paper, we assume the number of connected elements, i.e., $a$, is predefined.
@@ II-B System Model
where ${\bf{y}}^{b}$ and ${\bf{y}}^{c}$ represent the signals received by the BS’s antennas and the RDARS’s elements operating in the connection mode, respectively. The effective channel ${\bf{H}}$ is composed of two parts, i.e., ${\bf{H}}_{b}$ and ${\bf{H}}_{c}$. First, ${\bf{H}}_{b}$ represents the effective channel between users and the BS assisted with the RDARS’s elements in the reflection mode, similar to that of the conventional RIS-aided system, defined as ${\bf{H}}_{b}={\bf{H}}_{d}+{\bf{G}}^{H}({\bf{I}}\!-\!{\bf{A}}){\bf{\Phi}}{\bf{H}}_{r}$. On the other hand, ${\bf{H}}_{c}$ denotes the effective channel between the users and the elements working in the connection mode, i.e., ${\bf{H}}_{c}={\bf{A}}_{a}{\bf{H}}_{r}$, where ${\bf{A}}_{a}\in\mathbb{C}^{a\times N}$ stands for the channel index matrix of those components connected to the BS. Specifically, ${\bf{A}}_{a}$ is a submatrix of ${\bf{A}}$ consisting of non-zero rows of the mode selection matrix ${\bf{A}}$. Here, ${\bf{P}}$ denotes the transmit power matrix for $M$ users, i.e., ${\bf{P}}={\rm diag}([\sqrt{p_{1}},\sqrt{p_{2}},\dots,\sqrt{p_{M}}])$, with $p_{m}$ representing the transmit power for the $m$-th user. ${\bf{n}}_{b}$ and ${\bf{n}}_{c}$ represent the additive white Gaussian noise (AWGN) obeying the CSCG distributions with zero mean and covariance matrices ${\bm{\Lambda}}_{b}\!=\!{\sigma_{b}^{2}}{\bf{I}}_{N_{r}}$ and ${\bm{\Lambda}}_{c}\!=\!{\sigma_{c}^{2}}{\bf{I}}_{a}$, i.e., ${\bf{n}}_{b}\sim\mathcal{CN}({\bf{0}},{\sigma_{b}^{2}}{\bf{I}}_{N_{r}})$ and ${\bf{n}}_{c}\sim\mathcal{CN}({\bf{0}},{\sigma_{c}^{2}}{\bf{I}}_{a})$, respectively. ${\bm{\Lambda}}={\rm blkdiag}{({\bm{\Lambda}}_{b},{\bm{\Lambda}}_{c})}$ is the covariance matrix of the effective noise vector ${\bf{n}}$.
@@ II-C Problem Formulation
The average sum MSE minimization problem for the RDARS-aided MIMO system can be formulated as
@@
where (3b) denotes the unit-modulus constraints for the passive reflection coefficients at the RADAS, and (3c)(3d) imply the mode selection matrix ${\bf{A}}$ is diagonal and has diagonal value of 0 or 1, indicating whether the RDARS element is connected to the BS. The cardinality constraint (3e) limits the number of elements in the connection mode to $a$. Unfortunately, problem (P1) is essentially a mixed-integer optimization problem, which is intractable and NP-hard. Due to the cardinality and binary constraints, problem (P1) is more challenging than the conventional fully passive RIS-aided system design. To the best of our knowledge, there have been very few studies considering dynamically configuring the RIS elements to be active or not. Additionally, existing algorithms are also inapplicable to our works. In the following section, we propose two different algorithms to overcome the challenges arising from unit-modulus, binary, and cardinality constraints.
@@ III Proposed Inexact BCD-PDD Algorithm
By taking the first derivative w.r.t ${\bf{W}}$ to 0, the optimal receive beamforming matrix ${\bf{W}}^{\star}$ can be obtained as
@@
Substituting ${\bf{W}}^{\star}$ into the objective function $f_{\rm MSE}^{{\bf{A}},{\bf{W}},{\bm{\theta}}}$, we have
@@
where the matrix-valued quadratic function ${\bf{R}}_{b}({\bf{A}},{\bf{\Phi}})$ and ${\bf{R}}_{c}({\bf{A}})$ w.r.t. ${\bf{A}}$ and ${\bf{\Phi}}$ are respectively defined as
${\bf{R}}_{b}({\bf{A}},{\bf{\Phi}})={\bf{P}}^{H}{\bf{H}}_{r}^{H}({\bf{I}}\!-\!{\bf{A}}){\bf{\Phi}}^{H}{\bf{G}}{\bf{G}}^{H}{\bf{\Phi}}({\bf{I}}\!-\!{\bf{A}}){\bf{H}}_{r}{\bf{P}}$
$\quad-\!2\Re\{{\bf{P}}^{H}{\bf{H}}_{d}^{H}{\bf{G}}^{H}{\bf{\Phi}}({\bf{I}}\!-\!{\bf{A}}){\bf{H}}_{r}{\bf{P}}\}+{{\bf{P}}^{H}{\bf{H}}_{d}^{H}{\bf{H}}_{d}{\bf{P}}}$ (6)
${\bf{R}}_{c}({\bf{A}})={{\bf{P}}^{H}{\bf{H}}_{r}^{H}{\bf{A}}{\bf{H}}_{r}{\bf{P}}}$ (7)
The sum-MSE minization problem for jointly optimizing the mode selection matrix and the RDARS’s phase shifts is then formulated as:
@@ III-A Equivalent Transformation
where $f_{\rm MSE}^{{\bf{x}},{\bf{v}},{\bm{\theta}}}={\rm Tr}\left\{({\bf{I}}_{M}\!+\!\sigma_{b}^{-2}{\bf{R}}_{b}({\bf{A}},{\bf{\Phi}})\!+\!\sigma_{c}^{-2}{\bf{\bar{R}}}_{c}({\bf{\bar{A}}}))^{-1}\right\}$ and ${\bf{\bar{R}}}_{c}({\bf{\bar{A}}})$ is obtained by replacing ${\bf{\bar{A}}}$ with ${\bf{A}}$ in (7).
@@
By virtue of Lemma 1, we can establish the equivalence between problem (P2) and problem (P3). ∎
@@ III-B2 ${\bf{x}}$-subproblem
However, due to the indefiniteness of ${\bf{\Xi}}$, the box-constrained optimization problem is generally not a standard convex problem. To efficiently address the challenges, we divide the objective into the convex and concave parts and transform it into a box-constrained difference-of-convex (DC) problem. The convex-concave procedure (CCP) [22],[23] is then proposed to find a local optimum by replacing the concave terms with a convex upper bound.
@@ III-D Discussion
Finally, the inexact BCD-based PDD algorithm for solving the problem (P2) is summarized in Algorithm 1. The proposed algorithm is guaranteed to converge to a stationary solution via iterative update of the primal and dual variables, as well as the penalty factor [26]. The initial values of the dual variables and the penalty parameter should be selected carefully to ensure the penalty term remains sufficiently small. Additionally, instead of setting a maximum number of iterations, we choose the relative objective progress (RBP) condition as the termination criterion, i.e.,
@@ III-E Special Cases
In this subsection, we mainly explore two specific cases, i.e., the single-input single-output (SISO) case and the line-of-sight (LoS) case, under the single-user scenario. We aim to investigate the intuitive beamforming and the mode selection solution for the MSE minimization problem in these cases.
In the RDARS-aided single-user communication system, the average MSE minimization problem (P2) degenerates to the following maximization problem
@@
Then, the optimal closed-form solutions are derived to give insightful observations for the intuitive beamforming design and the mode selection optimzation in these two special cases.
III-E1 Single-antenna case
@@
For the RDARS-aided single-user SISO system, the selection matrix ${\bf{A}}$ can be simplified to select the first $a$ largest channels between the user and the RADRS, i.e., ${\mathcal{A}=\{c_{1},c_{2},...,c_{a}\}},$ where the indices $c_{i}$ are defined according to the rank ordering of the channel gain, i.e., $h_{r,c_{1}}^{2}\!\geq\!\dots\!\geq\!h_{r,c_{N}}^{2}$. The reflection coefficients are optimally solved by ${\bm{\theta}}^{\star}=e^{j(\angle{{{h}}_{d}}-\angle{{\bf{h}}_{r}}+\angle{{\bf{g}}})}$, where ${\bf{g}}$ denotes the reduced channel ${\bf{G}}$ in the SISO case.
@@ III-E2 LoS case
Assume the LoS BS-RDARS channel ${\bf{G}}=\kappa_{g}{\bf{a}}_{gr}{\bf{a}}_{gt}^{H}$ and the LoS RDARS-user channel ${\bf{h}}_{r}=\kappa_{r}{\bf{a}}_{r}$, then the selection matrix ${\bf{A}}$ can be chosen arbitrarily. The reflection coefficients can be optimally solved with ${\bm{\theta}}^{\star}=e^{j(\angle{\bf{{\tilde{a}}}_{r}}-\angle{{\bf{h}}_{d}^{H}{\bf{a}}_{t}})}.$, where ${\bf{\tilde{a}}}_{r}=\kappa_{g}{\bf{a}}_{gr}^{H}({\bf{I}}\!-\!{\bf{A}}){\rm diag}({\bf{h}}_{r})$.
@@ 11       $k=k+1$
12until  RBP termination criterion is met ;
@@ IV-B Beamforming Design
14until  $|f_{\rm MSE}^{app}({\mathcal{A}}_{k},{\bm{\theta}}_{k})-f_{\rm MSE}^{app}({\mathcal{A}}_{k\!-\!1},{\bm{\theta}}_{k\!-\!1})|\leq\epsilon$ ;
@@ V Simulations
In this section, we provide numerical simulations to evaluate the performance of all proposed algorithms for the RDARS-aided communication system. Unless otherwise stated, we set the basic system parameters as follows: $N_{r}=4$, $M=4$, $N=256$. Under a three-dimensional deployment setup as shown in Fig. 3, the coordinate of the BS is set as $(0,100,5)$m and the users are randomly distributed in a circle with a radius of $10$m and a center of $(0,0,1.5)$m. The location of RDARS are $(0,50,15)$m and $(0,100,15)$m, respectively. The path loss model at 3GPP [29] is considered, i.e., $\text{PL}(d)\!=\!\beta_{0}\Psi_{a}{(d)}^{-{\alpha}_{p}}$, where $d$ and ${\alpha}_{p}$ respectively denote the propagation distance and the path loss exponent. $\beta_{0}$ stands for the path loss at the reference distance 1m and $\Psi_{a}$ denotes the random shadowing impact subject to the Gaussian distribution with zero mean and a standard derivation of $\sigma$. We assume $\beta_{0}\!=\!-30$dB and $\Psi_{a}\sim\mathcal{N}(0,\sigma^{2})$ with $\sigma=5.8$dB in this simulation. In particular, we set ${\alpha}_{p}^{RB}=2.2$, ${\alpha}_{p}^{UR}=2.2$ and ${\alpha}_{p}^{UB}=3.5$ for the BS-RDRAS channel ${\bf H}_{RB}$, the RDRAS-user channels ${\bf H}_{UR}$ and BS-user channels ${\bf H}_{UB}$, respectively.
As for the small-scale fading, we assume the Rician fading channels, modeled as
@@
where ${\bf{H}}^{\rm LoS}$ and ${\bf{H}}^{\rm NLoS}$ represent the deterministic LoS and non-LoS (NLoS) components, respectively. The NLoS component ${\bf{H}}^{\rm NLoS}$ is assumed to be Rayleigh fading, where each entry of ${\bf{H}}^{\rm NLoS}$ subjects to the CSCG distribution with $\mathcal{CN}(0,1)$. The deterministic LoS component is determined by the antenna array response vectors, i.e., ${\bf{h}}_{UB,m}^{\rm LoS}={\bf{a}}_{B}(\psi_{UB,m})$, ${\bf{G}}^{\rm LoS}={\bf{a}}_{R}(\psi_{A},\vartheta_{A}){\bf{a}}_{B}(\psi_{RB})^{H}$ and ${\bf{h}}_{UR,m}^{\rm LoS}={\bf{a}}_{R}(\psi_{D,m},\vartheta_{D,m})$, where ${\bf{a}}_{B}$ and ${\bf{a}}_{R}$ denote the array response vectors at the BS and the RDARS, respectively. $\psi_{UB,m}$ and $\psi_{RB}$ are the angles of departure (AoDs) at the BS towards the $m$-th user and the RDARS, respectively. Furthermore, $\psi_{A}$ and $\vartheta_{A}$ denote the azimuth and elevation angles of arrival (AoAs) at the RDARS, respectively, whereas $\psi_{D}$ and $\vartheta_{D}$ stand for the AoDs. $\kappa_{t}$ in (42) represents the normalized Rician factor belonging to $[0,1]$. With $\kappa_{t}$ increasing from 0 to 1, the channel ${\bf{H}}$ at Rayleigh fading becomes the LoS channel. Specifically, we assume the channel between users and the BS ${\bf{H}}_{d}$ is assumed to be Rayleigh fading, while the RDARS-related channels ${\bf{H}}_{r}$ and ${\bf{G}}$ obey Rician fading with Rician factor $\kappa_{t}$ being 0.75 in this simulation unless otherwise stated. The noise powers $\sigma_{b}^{2}$ and $\sigma_{c}^{2}$ are given by $\sigma_{b}^{2}=\sigma_{c}^{2}=\sigma^{2}=-90$dBm. The average normalized MSE (ANMSE) denoted by ${\rm ANMSE}\triangleq{\rm Tr}({\bf MSE})/M$ is adopted as the performance metric. All simulation results are obtained by averaging over $300$ channel realizations.
@@
V-B ANMSE vs $P$.

### v2

@@ Abstract
publicationid: pubid: © 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ II-A RDARS Architecture
A RDARS composed of $N$ elements is employed, with $a$ elements connected to the BS and operating in the “connection mode”, while the remaining $N\!-\!a$ passive elements operate in the “reflection mode” to reflect the incident signal to favorable directions. Since each element can be dynamically switched between the connection and reflection modes, it offers an extra DoF to achieve selection gains. Such transformation can be realized via a switching network (SW) as illustrated in Fig. 1 111 Compared to conventional DAS and RIS, the additional hardware in the RDARS is the SW, which includes RF switches designed for dynamic mode switching. As we know, RF switches have been utilized in antenna selection techniques within the domain of hybrid beamforming for massive MIMO communications. They serve to reduce hardware costs by substituting for analog phase shifters or RF chains [18, 19, 20]. Analog RF switches offer significantly lower complexity and power consumption, which enhances the appeal of antenna selection [21, 22]. Although the cable connections and mode switches involved in the hardware implementation of RDARS will incur an acceptable hardware cost, the performance improvement makes the RDARS structure particularly appealing.. Flexible configurations are available for the number and locations of elements in the connection mode. In this paper, we assume the number of connected elements, i.e., $a$, is predefined.
@@ II-B System Model
where ${\bf{y}}^{b}$ and ${\bf{y}}^{c}$ represent the signals received by the BS’s antennas and the RDARS’s elements operating in the connection mode, respectively. The effective channel ${\bf{H}}$ is composed of two parts, i.e., ${\bf{H}}_{b}$ and ${\bf{H}}_{c}$. First, ${\bf{H}}_{b}$ represents the effective channel between users and the BS assisted with the RDARS’s elements in the reflection mode, similar to that of the conventional RIS-aided system, defined as ${\bf{H}}_{b}={\bf{H}}_{d}+{\bf{G}}^{H}({\bf{I}}\!-\!{\bf{A}}){\bf{\Phi}}{\bf{H}}_{r}$. On the other hand, ${\bf{H}}_{c}$ denotes the effective channel between the users and the elements working in the connection mode, i.e., ${\bf{H}}_{c}={\bf{A}}_{a}{\bf{H}}_{r}$, where ${\bf{A}}_{a}\in\mathbb{C}^{a\times N}$ stands for the channel index matrix of those components connected to the BS. Specifically, ${\bf{A}}_{a}$ is a submatrix of ${\bf{A}}$ consisting of non-zero rows of the mode selection matrix ${\bf{A}}$. Here, ${\bf{P}}$ denotes the transmit power matrix for $M$ users, i.e., ${\bf{P}}={\rm diag}([\sqrt{p_{1}},\sqrt{p_{2}},\dots,\sqrt{p_{M}}])$, with $p_{m}$ representing the transmit power for the $m$-th user. ${\bf{n}}_{b}$ and ${\bf{n}}_{c}$ represent the additive white Gaussian noise (AWGN) obeying the CSCG distributions with zero mean and covariance matrices ${\bm{\Lambda}}_{b}\!=\!{\sigma_{b}^{2}}{\bf{I}}_{N_{r}}$ and ${\bm{\Lambda}}_{c}\!=\!{\sigma_{c}^{2}}{\bf{I}}_{a}$, i.e., ${\bf{n}}_{b}\sim\mathcal{CN}({\bf{0}},{\sigma_{b}^{2}}{\bf{I}}_{N_{r}})$ and ${\bf{n}}_{c}\sim\mathcal{CN}({\bf{0}},{\sigma_{c}^{2}}{\bf{I}}_{a})$, respectively. ${\bm{\Lambda}}={\rm blkdiag}{({\bm{\Lambda}}_{b},{\bm{\Lambda}}_{c})}$ is the covariance matrix of the effective noise vector ${\bf{n}}$. For simplicity, we assume ${\sigma_{b}^{2}}={\sigma_{c}^{2}}={\sigma^{2}}$. Thus, the covariance matrix of the effective noise vector turns into ${\bm{\Lambda}}={\sigma^{2}}{\bf{I}}$.
@@ II-C Problem Formulation
The average sum MSE minimization problem for the RDARS-aided MIMO system can be formulated as444 The proposed algorithms in this paper can be easily extended to the general case with the inequality constraint $||{\rm diag}({\bf{A}})||_{0}\leq a$. Additionally, we conclude that this general problem, subject to the constraint $||{\rm diag}({\bf{A}})||_{0}\leq a$, is equivalent to the problem with the equality constraint $||{\rm diag}({\bf{A}})||_{0}=a$.
@@
where (3b) denotes the unit-modulus constraints for the passive reflection coefficients at the RADAS, and (3c)(3d) imply the mode selection matrix ${\bf{A}}$ is diagonal and has a diagonal value of 0 or 1, indicating whether the RDARS element is connected to the BS. The cardinality constraint (3e) limits the number of elements in the connection mode to $a$. Unfortunately, problem (P1) is essentially a mixed-integer optimization problem, which is intractable and NP-hard. Due to the cardinality and binary constraints, problem (P1) is more challenging than the conventional fully passive RIS-aided system design. To the best of our knowledge, there have been very few studies considering dynamically configuring the RIS elements to be active or not. Additionally, existing algorithms are also inapplicable to our works. In the following section, we propose two different algorithms to overcome the challenges arising from unit-modulus, binary, and cardinality constraints.
@@ III Proposed Inexact BCD-PDD Algorithm
Since the objective function $f_{\rm MSE}^{{\bf{A}},{\bf{W}},{\bf{\Phi}}}$ is convex w.r.t. the receive beamforming matrix $\bf{W}$. The optimal solution can be obtained with closed form by taking the first derivative to 0, i.e.,
@@
Substituting the optimal receive beamforming matrix $\bf{W}^{\star}$ into the objective function in problem (P1), the newly formulated objective function with two optimization variables is derived as
@@
where ${\bf{R}}_{b}({\bf{A}},{\bf{\Phi}})={{\bf{P}}^{H}{\bf{H}}_{b}^{H}{\bf{H}}_{b}{\bf{P}}}$ and ${\bf{R}}_{c}({\bf{A}})={{\bf{P}}^{H}{\bf{H}}_{c}^{H}{\bf{H}}_{c}{\bf{P}}}$ denotes the received signal covariance matrices related to the effective channels ${\bf{H}}_{b}$ and ${\bf{H}}_{c}$, respectively.
The sum-MSE minimization problem for jointly optimizing the mode selection matrix and the RDARS’s phase shifts is then formulated as:
@@ III-A Equivalent Transformation
where $f_{\rm MSE}^{{\bf{x}},{\bf{v}},{\bm{\theta}}}={\rm Tr}\left\{({\bf{I}}_{M}\!+\!\sigma_{b}^{-2}{\bf{R}}_{b}({\bf{A}},{\bf{\Phi}})\!+\!\sigma_{c}^{-2}{\bf{\bar{R}}}_{c}({\bf{\bar{A}}}))^{-1}\right\}$ and ${\bf{\bar{R}}}_{c}({\bf{\bar{A}}})$ is obtained by replacing ${\bf{\bar{A}}}$ with ${\bf{A}}$ in ${\bf{R}}_{c}({\bf{A}})$.
@@
By virtue of Lemma 1, it is readily inferred that ${\bf{x}}\in\{0,1\}^{N}$ and ${\bf{x}}={\bf{v}}$ from the $l_{2}$ box non-separable constraints, i.e., ${\bf{0}}\leq{\bf{x}}\leq{\bf{1}},||2{\bf{v}}\!-\!{\bf{1}}||_{2}^{2}\leq N,\left<2{\bf{x}}\!-\!{\bf{1}},2{\bf{v}}\!-\!{\bf{1}}\right>=N$. Then, we can rewrite problem (P2) in an equivalent form as shown in problem (P3). Note that the newly formed constraint $(2{\bf{x}}\!-\!1)^{T}(2{\bf{v}}\!-\!1)=N$ is referred to as the complementarity constraint and it always holds that $(2{\bf{x}}\!-\!1)^{T}(2{\bf{v}}\!-\!1)\leq N$ for any feasible ${\bf{x}}$ and ${\bf{v}}$.
@@ III-B2 ${\bf{x}}$-subproblem
However, due to the indefiniteness of ${\bf{\Xi}}$, the box-constrained optimization problem is generally not a standard convex problem. To efficiently address the challenges, we divide the objective into the convex and concave parts and transform it into a box-constrained difference-of-convex (DC) problem. The convex-concave procedure (CCP) is then proposed to find a local optimum by replacing the concave terms with a convex upper bound.
@@ III-D Discussion
Finally, the inexact BCD-based PDD algorithm for solving the problem (P2) is summarized in Algorithm 1. The proposed algorithm is guaranteed to converge to a stationary solution via an iterative update of the primal and dual variables, as well as the penalty factor [31]. The initial values of the dual variables and the penalty parameter should be selected carefully to ensure the penalty term remains sufficiently small. Additionally, instead of setting a maximum number of iterations, we choose the relative objective progress (RBP) condition as the termination criterion, i.e.,
@@ III-E Special Cases
In this subsection, we focus primarily on the RDARS-aided single-user system and explore two specific cases, i.e., the single-antenna BS case and the line-of-sight (LoS) case. We aim to investigate the intuitive beamforming and the mode selection solution for the MSE minimization problem in these cases.
In the RDARS-aided one single-antenna user communication system, the average MSE minimization problem (P2) degenerates to the following maximization problem
@@
where ${\bf{h}}_{d}$ and ${\bf{h}}_{r}$ denote the reduced channels from the users to the BS and the RDARS, i.e., ${\bf{H}}_{d}$ and ${\bf{H}}_{r}$, in the RDARS-aided single-user scenario, respectively. Then, the optimal closed-form solutions are derived to give insightful observations for the intuitive beamforming design and the mode selection optimization in these two special cases.
III-E1 Single-antenna BS case
@@
For the RDARS-aided one single-antenna user system with single-antenna BS, the selection matrix ${\bf{A}}$ can be simplified to select the first $a$ largest channels between the user and the RADRS, i.e., ${\mathcal{A}=\{c_{1},c_{2},...,c_{a}\}},$ where the indices $c_{i}$ are defined according to the rank ordering of the channel gain, i.e., $h_{r,c_{1}}^{2}\!\geq\!\dots\!\geq\!h_{r,c_{N}}^{2}$. The reflection coefficients are optimally solved by ${\bm{\theta}}^{\star}=e^{j(\angle{{{h}}_{d}}-\angle{{\bf{h}}_{r}}+\angle{{\bf{g}}})}$, where ${\bf{g}}$ denotes the reduced channel ${\bf{G}}$ in the single-antenna BS case.
@@ III-E2 LoS case
Assume the LoS BS-RDARS channel ${\bf{G}}=\kappa_{g}{\bf{a}}_{gr}{\bf{a}}_{gt}^{H}$, where ${\bf{a}}_{gt}$ denotes the transmit steering vector of the BS antennas and ${\bf{a}}_{gr}$ represents the receive steering vector of the RDARS elements. Similarly, the LoS RDARS-user channel is assumed to be ${\bf{h}}_{r}=\kappa_{r}{\bf{a}}_{r}$ with the RDARS transmit steering vector ${\bf{a}}_{r}$. $\kappa_{g}$ and $\kappa_{r}$ denote the large-scale channel coefficients for BS-RDARS and RDARS-user channels, respectively. Then, the selection matrix ${\bf{A}}$ can be chosen arbitrarily and the reflection coefficients can be optimally solved with ${\bm{\theta}}^{\star}=e^{j(\angle{\bf{{\tilde{a}}}_{r}}-\angle{{\bf{h}}_{d}^{H}{\bf{a}}_{t}})}.$, where ${\bf{\tilde{a}}}_{r}=\kappa_{g}{\bf{a}}_{gr}^{H}({\bf{I}}\!-\!{\bf{A}}){\rm diag}({\bf{h}}_{r})$.
@@ 12       $k=k+1$
13until  RBP termination criterion is met ;
@@ IV-B Beamforming Design
15until  $|f_{\rm MSE}^{app}({\mathcal{A}}_{k},{\bm{\theta}}_{k})-f_{\rm MSE}^{app}({\mathcal{A}}_{k\!-\!1},{\bm{\theta}}_{k\!-\!1})|\leq\epsilon$ ;
@@ V Simulations
In this section, we provide numerical simulations to evaluate the performance of all proposed algorithms for the RDARS-aided communication system. Unless otherwise stated, we set the basic system parameters as follows: $N_{r}=4$, $M=4$, $N=256$. Under a three-dimensional deployment setup as shown in Fig. 3, the coordinate of the BS is set as $(0,200,5)$m and the users are randomly distributed in a circle with a radius of $10$m and a center of $(0,0,1.5)$m. The location of RDARS is $(30,d_{R},15)$m, where $d_{R}$ denotes the distance of the RDARS along the y-axis. The path loss model at 3GPP [34] is considered, i.e., $\text{PL}(d)\!=\!\beta_{0}\Psi_{a}{(d)}^{-{\alpha}_{p}}$, where $d$ and ${\alpha}_{p}$ respectively denote the propagation distance and the path loss exponent. $\beta_{0}$ stands for the path loss at the reference distance 1m and $\Psi_{a}$ denotes the random shadowing impact subject to the Gaussian distribution with zero mean and a standard derivation of $\tilde{\sigma}$. We assume $\beta_{0}\!=\!-30$dB and $\Psi_{a}\sim\mathcal{N}(0,{\tilde{\sigma}}^{2})$ with ${\tilde{\sigma}}=5.8$dB in this simulation. In particular, we set ${\alpha}_{p}^{RB}=2.2$, ${\alpha}_{p}^{UR}=2.2$ and ${\alpha}_{p}^{UB}=3.5$ for the BS-RDRAS channel ${\bf H}_{RB}$, the RDRAS-user channels ${\bf H}_{UR}$ and BS-user channels ${\bf H}_{UB}$, respectively.
As for the small-scale fading, we assume the spatially correlated Rician fading channels, modeled as
@@
where ${\bf{H}}^{\rm LoS}$ and ${\bf{H}}^{\rm NLoS}$ represent the deterministic LoS and spatial non-LoS (NLoS) components, respectively. $\kappa_{t}$ in (40) represents the normalized Rician factor belonging to $[0,1]$. With $\kappa_{t}$ increasing from 0 to 1, the channel ${\bf{H}}$ transitions from a Rayleigh fading channel to an LoS channel.
The deterministic LoS component is determined by the antenna array response vectors, i.e., ${\bf{h}}_{UB,m}^{\rm LoS}={\bf{a}}_{B}(\psi_{UB,m})$, ${\bf{G}}^{\rm LoS}={\bf{a}}_{R}(\psi_{A},\vartheta_{A}){\bf{a}}_{B}(\psi_{RB})^{H}$ and ${\bf{h}}_{UR,m}^{\rm LoS}={\bf{a}}_{R}(\psi_{D,m},\vartheta_{D,m})$, where ${\bf{a}}_{B}$ and ${\bf{a}}_{R}$ denote the array response vectors at the BS and the RDARS, respectively. $\psi_{UB,m}$ and $\psi_{RB}$ are the angles of departure (AoDs) at the BS towards the $m$-th user and the RDARS, respectively. Furthermore, $\psi_{A}$ and $\vartheta_{A}$ denote the azimuth and elevation angles of arrival (AoAs) at the RDARS, respectively, whereas $\psi_{D}$ and $\vartheta_{D}$ stand for the AoDs.
The NLoS component ${\bf{H}}^{\rm NLoS}$ is assumed to be correlated Rayleigh fading, modeled as
${\bf{H}}^{\rm NLoS}={\bf{R}}^{\frac{1}{2}}{\bf{W}}{\bf{T}}^{\frac{1}{2}}$ (41)
where ${\bf{W}}$ denotes the random components of each channel with independent and identically distributed (i.i.d.) zero-mean unit covariance complex Gaussian random elements. ${\bf{R}}$ and ${\bf{T}}$ are deterministic nonnegative semi-definite receive and transmit spatial correlation matrices, respectively. Specifically, ${\bf{H}}^{\rm NLoS}$ becomes the conventional Rayleigh fading channel when ${\bf{R}}={\bf{T}}={\bf{I}}$. We adopt the exponential correlation model to characterize the spatial correlation among the antennas [35, 36]. For the ULA topology at the BS, the (i,j)-th element of correlation matrix $R_{B}$ is given as $R_{B}(i,j)=\beta_{R_{B}}^{|i-j|}$. Regarding the spatial transmit and receive correlation matrices ${T}_{R}$ and ${R}_{R}$ at the RDARS, we apply the Kronecker product of the vertical correlation matrix ${R}_{v}$ and the horizontal correlation matrix ${R}_{h}$ to the correlation model and assume ${T}_{R}={R}_{R}$, i.e., ${T}_{R}={R}_{R}={R}_{v}\otimes{R}_{h}$ [37, 38]. Using the exponential correlation model, the element in ${R}_{v}$ and ${R}_{h}$ are given as $R_{v}(i,j)=\beta_{R_{v}}^{|i-j|}$ and $R_{h}(i,j)=\beta_{R_{h}}^{|i-j|}$, respectively. $\{\beta_{R_{B}},\beta_{R_{v}},\beta_{R_{h}}\}$ denotes the correlation coefficients between any two adjacent antennas or elements at the BS and the RDARS. The correlation coefficients at the RDARS are set as $\beta_{R_{v}}=\beta_{R_{h}}=\beta_{c}=0.5$. In contrast, the correlation coefficient at the BS is set as $\beta_{R_{B}}=0$ unless otherwise stated.
Specifically, we assume the channel between users and the BS ${\bf{H}}_{d}$ is assumed to be Rayleigh fading with Rician factor $\kappa_{t}$ being 0, while the RDARS-related channels ${\bf{H}}_{r}$ and ${\bf{G}}$ obey Rician fading with Rician factor $\kappa_{t}$ being 0.75 in this simulation unless otherwise stated. The noise powers $\sigma_{b}^{2}$ and $\sigma_{c}^{2}$ are given by $\sigma_{b}^{2}=\sigma_{c}^{2}=\sigma^{2}=-90$dBm. The average normalized MSE (ANMSE) denoted by ${\rm ANMSE}\triangleq{\rm Tr}({\bf MSE})/M$ is adopted as the performance metric. All simulation results are obtained by averaging over $300$ channel realizations.
@@
Figure 5: Performance-complexity trade-off between IBCD-PDD and GS-AO methods.
@@
V-B Performance-Complexity Trade-off.
Fig. 5 illustrates the performance-complexity trade-off between the proposed IBCD-PDD and GS-AO methods. The IBCD-PDD method demonstrates superior ANMSE performance over the GS-AO method as the scale of RDARS increases. However, the CPU time required by the IBCD-PDD method increases sharply as the number of RDARS elements $N$ increases, which is always higher than that of the GS-AO method. This pronounced difference is because the computational complexity of the IBCD-PDD method scales with $N$ to the 3.5th power, in contrast to the quadratic scaling of the GS-AO method. Therefore, the superior performance of the IBCD-PDD comes at the cost of increased computational complexity when compared to the GS-AO method. This relationship underscores the necessity for a judicious selection of the method based on the specific requirements and constraints of the application at hand, such as the scale of the RDARS system and the available computational resources.
V-C ANMSE vs $P$.